# Circumbinary disks in post common envelope binary systems with compact objects

### v1

@@ 6 Summary and discussion
In this study, we analyzed the population synthesis data of massive binaries from Grichener (2023), as described in Section 2, to explore the formation of post-CEE CBDs in systems that evolve into DCO binaries.
@@
When an NS or BH engages in a CEE phase with a giant star, it spirals inside the envelope and ejects it. Some former envelope mass might become a close equatorial circumbinary gas, i.e., a CBD around the exposed giant’s core and the compact object. Since this CBD formation process is challenging to simulate, we parametrized it with a toy model, as described in Section 3.1. We assume that the leftover envelope gas has a specific angular momentum that is a fraction $\beta$ of the synchronized envelope at the onset of CEE (equation 2). If this specific angular momentum is larger than of a Keplerian disk around the core-compact object binary system, than the left-over envelope forms a CBD (equation 3); this is quantified in equation 4. In this exploratory study, we analyzed results for values of $\beta$ that lead half of NS-NS, BH-BH and NS-BH binaries to form a CBD, as inferred from Figure 1.
@@
CBDs will form most easily in BH-BH binary progenitors (Figure 3 in Section 4). Within our toy model, the CBD formation mostly depends on the depth the compact object manages to dive to during the CEE phase (Left columns of Figures 2 - 4), which is the smallest for BH-BH binaries where the secondary stars are the most massive. For DCO binaries that are not expected to merge within Hubble time, forming a post-CEE CBD is less likely due to the relatively large orbital separation at the end of CEE. In the case of NS-BH binaries, however, the larger mass ratio between the BH and the progenitor of the NS facilitates the formation of the CBD.
@@
Our results affect the predicted formation and merger rates of DCO binaries. Angular momentum transfer between the CBD and the binary system dictates the final post-CEE orbital evolution (e.g., Tuna & Metzger 2023; Valli et al. 2024; Vetter et al. 2024). We found that the number of DCO mergers will likely decrease independently of whether the binary widens or contracts due to interactions with the CBD. In cases where the orbital separation grows, the DCOs in the binary might drift so far apart they may no longer be able to merge within Hubble time. If the CBD-binary interaction decreases the orbit, on the other hand, in some of the systems that were supposed to evolve into DCO mergers, the compact object and the core will merge shortly after the CEE. This likely results in luminous supernova-like transients when the core is composed of Helium or long gamma-ray bursts for more evolved cores. These transients are termed common envelope jets supernovae (CEJSNe; Soker & Gilkis 2018). Cases where the compact objects does not merge with the core are CEJSN impostors (e.g., Gilkis et al. 2019).
@@
The possible contraction of the orbital separation might also result in shorter delay time distributions for DCO mergers, with potential implications for r-process nucleosynthesis, i.e., allows NS-NS merger to synthesis r-process elements at earlier galactic evolution times. Since most merging DCO binaries are expected to go through a CEE phase of a compact object and a giant star (middle panel of Figure 6), we expect that CBD interactions would have a non-negligible effect in lowering the merger rates of DCO binaries. While quantifying the interaction between post-CEE CBDs and the binary systems they encompass to find the resultant orbital evolution and determining the lowering factor is a challenging task that is beyond the scope of this paper, we conclude it is important to keep in mind that CBD interaction might affect the rates of observables predicted by binary population synthesis simulations.
@@
Different choices of parameters and prescriptions in our population synthesis model can influence the results we obtain. Metallicity has shown to have a non-negligible impact on the formation and merger rates of DCO binaries, with a greater influence on BH-BH formation that is expected to be dominant at lower metallicities (e.g. Mapelli et al. 2017; Giacobbo & Mapelli 2018; Neijssel et al. 2019; van Son et al. 2024). We found that the formation and merger rates of BH-BH binaries grows significantly for $Z=0.01Z_{\rm\odot}$ with respect to solar metalicity, with a much smaller effect on NS-NS and NS-BH binaries. Most of the merging BH-BH systems still evolve through CEE and most non-merging BH-BH binaries cannot form CBDs, implying our main conclusions from this study hold, and that we can predict fewer detected gravitational wave events with Ligo-Virgo-Kagra (LVK, e.g., Abbott et al. 2018) compared to the expectations of population synthesis studies that neglect CBD formation. Finding the exact reduction factor would require simulations with a metallicity-dependent star formation history and parametrization of the CBD-binary interaction.
@@
Mass loss through stellar winds significantly influences stellar evolution and the compact objects that massive stars form (e.g., Smith 2014). Since compas does not account for wind accretion or wind interactions with the companion (Riley et al., 2022), we investigated the effect of conservative mass loss on CBD formation by running simulations accounting for accretion of stellar winds by adding to the accretor the mass lost by the donor. We found that wind accretion does not significantly impact our results. This finding reinforces our conclusion from Section 3.2 that the primary factor determining the ratio of the specific angular momentum of the CBD to the synchronized envelope-specific angular momentum at CEE onset, and consequently $\beta$, is the spiraling-in factor $\eta_{\rm SIF}\equiv a_{\rm f}/R_{\ast,i}$. Other uncertainties related to binary evolution, such as the representation of CEE, tidal evolution, and natal kick prescriptions, might affect the results of our study in non-trivial ways and will be a subject of future studies.
@@ ACKNOWLEDGMENTS
We thank Ruggero Valli and Semih Tuna for the helpful discussions. AG acknowledges support from the Miriam and Aaron Gutwirth Fellowship, the Steward Observatory Fellowship in Theoretical and Computational Astrophysics, the IAU-Gruber Fellowship, and the CHE Fellowship. Simulations in this paper used the compas rapid binary population synthesis code (version 02.31.06), which is freely available at http://github.com/TeamCOMPAS/COMPAS.
@@
The data underlying this article is publicly available at doi:10.5281/zenodo.11237179

### v2

@@ VI Summary and discussion
In this study, we analyzed the population synthesis data of massive binaries from Grichener (2023), as described in Section II, to explore the formation of post-CEE CBDs in systems that evolve into DCO binaries.
@@
When an NS or BH engages in a CEE phase with a giant star, it spirals inside the envelope and ejects it. Some former envelope mass might become a close equatorial circumbinary gas, i.e., a CBD around the exposed giant’s core and the compact object. Since this CBD formation process is challenging to simulate, we parametrized it with a toy model, as described in Section III.1. We assume that the leftover envelope gas has a specific angular momentum that is a fraction $\beta$ of the total specific angular momentum at the onset of RLOF (equation 4). If this specific angular momentum is larger than that of a Keplerian disk around the CBD’s inner radius, then the left-over envelope forms a CBD (equation 5). In this exploratory study, we analyzed results for values of $\beta$ that lead half of NS-NS, BH-BH, and NS-BH binaries to form a CBD, as inferred from Figure 1.
@@
CBDs will form least easily in NS-NS binary progenitors (Figure 2 in Section IV). For DCO binaries that are not expected to merge within Hubble time, forming a post-CEE CBD is less likely due to the relatively large orbital separation at the end of CEE. In the case of NS-BH binaries, however, the larger mass ratio between the BH and the progenitor of the NS facilitates the formation of the CBD.
@@
Our results affect the predicted formation and merger rates of DCO binaries. Angular momentum transfer between the CBD and the binary system dictates the final post-CEE orbital evolution (e.g., Tuna and Metzger 2023; Siwek et al. 2023; Valli et al. 2024; Wei et al. 2024). We found that the number of NS-NS and BH-BH mergers will likely decrease independently of whether the binary widens or contracts due to interactions with the CBD. In cases where the orbital separation grows, the DCOs in the binary might drift so far apart that they may no longer be able to merge within Hubble time. If the CBD-binary interaction decreases the orbit, on the other hand, in some of the systems that were supposed to evolve into DCO mergers, the compact object and the core will merge shortly after the CEE. This likely results in luminous supernova-like transients when the core is composed of Helium or long gamma-ray bursts for more evolved cores. These transients are CEJSNe. We note that without accounting for CBD interaction, there is a negligible number of CEJSN that occur from a merger shortly after CEE. Therefore, a scenario where a CBD widens the orbit preventing a CEJSN and leading to a DCO mergers instead, will not increase the number of merging DCOs.
@@
The possible contraction of the orbital separation might also result in shorter delay time distributions for DCO mergers, with potential implications for r-process nucleosynthesis, i.e., allowing NS-NS mergers to produce r-process elements at earlier stages of galactic evolution. Since we found most merging DCO binaries are expected to go through a CEE phase of a compact object and a giant star (middle panel of Figure 6), we expect that CBD interactions would have a non-negligible effect in lowering the merger rates of DCO binaries. While quantifying the interaction between post-CEE CBDs and the binary systems they encompass to find the resultant orbital evolution and determining the lowering factor is a challenging task that is beyond the scope of this paper, we conclude it is important to keep in mind that CBD interaction might affect the rates of observables predicted by binary population synthesis simulations.
@@
Different choices of parameters and prescriptions in our population synthesis model can influence the results we obtain. Metallicity has been shown to have a non-negligible impact on the formation and merger rates of DCO binaries, with a greater influence on BH-BH formation that is expected to be dominant at lower metallicities (e.g. Mapelli et al. 2017; Giacobbo and Mapelli 2018; Neijssel et al. 2019; van Son et al. 2024; but see Bavera et al. 2023; Lalleman et al. 2025 for another view). We found that the formation and merger rates of BH-BH binaries grow significantly for $Z=0.01Z_{\rm\odot}$ with respect to solar metalicity (Fig. 7), with a much smaller effect on NS-NS and NS-BH binaries. A large fraction of the merging BH-BH systems ($\simeq 67\%$) still evolve through CEE and most non-merging BH-BH binaries go through stable mass transfer and hence cannot form CBDs, implying our main conclusions from this study hold, and that we can predict fewer detected gravitational wave events with Ligo-Virgo-Kagra (LVK, e.g., Abbott et al. 2018) compared to the expectations of population synthesis studies that neglect CBD formation. Finding the exact reduction factor would require simulations with a metallicity-dependent star formation history, selection effects, and parametrization of the CBD-binary interaction.
@@
Mass loss through stellar winds significantly influences stellar evolution and the compact objects that massive stars form (e.g., Smith 2014). Since compas does not account for wind accretion or wind interactions with the companion (Riley et al., 2022), we investigated the effect of conservative mass loss on CBD formation by running simulations accounting for accretion of stellar winds by adding to the accretor the mass lost by the donor. We found that wind accretion does not significantly impact our results. Other uncertainties related to binary evolution, such as the representation of CEE, tidal evolution, and natal kick prescriptions, might affect the results of our study in non-trivial ways and will be a subject of future studies.
@@ ACKNOWLEDGMENTS
We thank Ruggero Valli and Semih Tuna for the helpful discussions, and an anonymous referee for valuable suggestions and comments. AG acknowledges support from the Miriam and Aaron Gutwirth Fellowship, the Steward Observatory Fellowship in Theoretical and Computational Astrophysics, the IAU-Gruber Fellowship, and the CHE Fellowship. Simulations in this paper used the compas rapid binary population synthesis code (version 02.31.06), which is freely available at http://github.com/TeamCOMPAS/COMPAS.
@@
This manuscript made use of data generated by Grichener (2023), publicly available at Zenodo: doi:10.5281. In particular, we re-analyzed the population synthesis data from the folders $\rm alpha\_1$ $\rm Z\_solar$ and $\rm alpha\_1$ $\rm 0.01Z\_solar$.